OpenStudy (anonymous):

MEDALS!!!!!!!!1Please help I'm so freakin confused!!!!! Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses.

OpenStudy (anonymous):

\[f(x)=\frac{ x+a }{ b }\] and \[g(x)=cx-d\]

OpenStudy (anonymous):

What are the numbers that are supposed to be in place of the letters for a, B, and C

OpenStudy (anonymous):

you could use a= 5 and x= 5 and b= 2 gx = 6 and c= 4 and d=9

OpenStudy (anonymous):

Thanks @caroloub